OpenStudy (anonymous):

What is the quantity of x plus 16 all over 3 = 3x?

OpenStudy (anonymous):

Is this correct? \[\frac{ x + 16 }{ 3 } = 3x\]

OpenStudy (anonymous):

if so, first you multiply both sides by 3 (to get the 3 out of the denominator) so you end up with 3x + 48 = 9x

OpenStudy (anonymous):

then subtract 3x from both sides, so you get 48 = 6x

OpenStudy (anonymous):

then divide both sides by 6, so you get x = 8

OpenStudy (anonymous):

X=2

OpenStudy (anonymous):

any other Questions?